# Energy landscape analysis based on the Ising model: Tutorial review

### v1

@@ 5 Applications
In our first application, we studied energy landscapes of fMRI data obtained from healthy humans undergoing a visual bistable perception task [6]. We used $N=7$ biologically informed locations in the brain, called regions of interest (ROI), as variables for constructing the PMEM. We constructed a single energy landscape from the fMRI data during the task. For the data pooled across the participants, we found ten local maxima and three heuristically determined major local maxima. In general, by inspecting the activity pattern at each major local minimum, one can see which ROI is relatively active (i.e., $\sigma_{i}(t)=1$) or inactive (i.e., $\sigma_{i}(t)=-1$), based on which one can biologically characterize the major local minima. One major state was an aggregation of five local minima and their basins, and we named this major state the frontal-area state because the ROIs in the frontal area of the brain were active at the major local minimum defining this major state. Likewise, the visual-area state was composed of two other local minima and their basins. The visual-area state is complementary to the frontal-area state in that the ROIs active in the former are inactive in the latter and vice versa. The third major state, named the intermediate state, contained three local minima. These three major states remained the same for the PMEM separately estimated for the data from each single participant. In the bistable visual perception task, the participants were asked to watch a set of moving dots on the screen, which allows bistable perception (i.e., either of the two directions of the apparently moving sphere composed of dots). The participants reported whenever their subjective perception spontaneously switched from one direction to the other. The main result was that participants with a larger basin size of the frontal-area state tended to have a shorter duration of the percept (i.e., how long the participant perceives one direction of motion of the visual stimulus before it subjectively switches to the other direction). In contrast, a larger basin size of the visual-area state tended to yield a longer duration of the percept. In this manner, we found an association between behavior (i.e., the median duration of the percept) and brain dynamics revealed by ELA (i.e., basin sizes), as schematically shown in Fig. 3(g).
@@ 6 Outlook
In section 3.1, we introduced a phase diagram analysis, which enables us to qualitatively tell properties of the given data set. If the system generating the observed time series data is changing on a slow timescale, which may be due to the environmental change (i.e., gradual pollution of an ecosystem) or the internal changes of the system (e.g., aging of the participant), the energy landscape of the system may transit from one phase to another (e.g, from the paramagnetic to spin-glass phase) on a slow timescale. A related challenge is to find early warning signals for large regime shifts of the system, which may occur as a result of bifurcations of the dynamical system underlying the data or for different reasons. Popular early warning signals include the sample variance and the decay rate of the sample autocorrelation function obtained from the observed time series, which would increase as the system approaches a bifurcation point [103, 104, 105, 106, 107]. A large regime shift that one wants to anticipate may correspond to a mass extinction of species in an ecosystem or a rapid progression of a disease in a human individual. From an energy landscape perspective, these methods usually assume that a large regime shift occurs when a local minimum in which the system’s state is currently sitting is lost via the diminishment of the energy barrier between the currently visited basin and a neighboring basin; see Fig. 1 in [103], Figs. 2 and 3 in [104], and Box 1 in [106] for schematics. Then, the regime shift corresponds to the movement of the system’s state from the current basin to the neighboring basin. Cross-fertilization of ELA and early warning signal research may be interesting.
@@ Acknowledgements
We acknowledge Pitambar Khanra for technical information. T.W. and N.M. acknowledge support from the Japan Science and Technology Agency (JST) Moonshot R&D (under grant no. JPMJMS2021). T.W. also acknowledges support from Grant-in-aid for Research Activity from Japan Society for Promotion of Sciences (23H04217), JST FOREST Program (24012854), Institute of AI and Beyond of UTokyo, and Cross-ministerial Strategic Innovation Promotion Program (SIP) on “Integrated Health Care System” (JPJ012425). N.M. also acknowledges support from National Institute of General Medical Sciences (under grant no. R01GM148973), the National Science Foundation (under grant no. 2204936), and JSPS KAKENHI (under grant nos. JP 23H03414, 24K14840, and W24K030130).

### v2

@@ 5 Applications
In our first application, we studied energy landscapes of fMRI data obtained from healthy humans undergoing a visual bistable perception task [6]. We used $N=7$ biologically informed locations in the brain, called regions of interest (ROI), as variables for constructing the PMEM. We constructed a single energy landscape from the fMRI data during the task. For the data pooled across the participants, we found ten local maxima and three heuristically determined major local maxima. In general, by inspecting the activity pattern at each major local minimum, one can see which ROI is relatively active (i.e., $\sigma_{i}(t)=1$) or inactive (i.e., $\sigma_{i}(t)=-1$), based on which one can biologically characterize the major local minima. One major state was an aggregation of five local minima and their basins, and we named this major state the frontal-area state because the ROIs in the frontal area of the brain were active at the major local minimum defining this major state. Likewise, the visual-area state was composed of two other local minima and their basins. The visual-area state is complementary to the frontal-area state in that the ROIs active in the former are inactive in the latter and vice versa. The third major state, named the intermediate state, contained three local minima. These three major states remained the same for the PMEM separately estimated for the data from each single participant. In the bistable visual perception task, the participants were asked to watch a set of moving dots on the screen, which allows bistable perception (i.e., either of the two directions of the apparently moving sphere composed of dots). The participants reported whenever their subjective perception spontaneously switched from one direction to the other. The main result was that participants with a larger basin size of the frontal-area state tended to have a shorter duration of the percept (i.e., how long the participant perceives one direction of motion of the visual stimulus before it subjectively switches to the other direction). In contrast, a larger basin size of the visual-area state tended to yield a longer duration of the percept. In this manner, we found an association between behavior (i.e., the median duration of the percept) and brain dynamics revealed by ELA (i.e., basin sizes), as schematically shown in Fig 3(g).
@@ 6 Outlook
In Section 3.1, we introduced a phase diagram analysis, which enables us to qualitatively tell properties of the given data set. If the system generating the observed time series data is changing on a slow timescale, which may be due to the environmental change (i.e., gradual pollution of an ecosystem) or the internal changes of the system (e.g., aging of the participant), the energy landscape of the system may transit from one phase to another (e.g, from the paramagnetic to spin-glass phase) on a slow timescale. A related challenge is to find early warning signals for large regime shifts of the system, which may occur as a result of bifurcations of the dynamical system underlying the data or for different reasons. Popular early warning signals include the sample variance and the decay rate of the sample autocorrelation function obtained from the observed time series, which would increase as the system approaches a bifurcation point [103, 104, 105, 106, 107]. A large regime shift that one wants to anticipate may correspond to a mass extinction of species in an ecosystem or a rapid progression of a disease in a human individual. From an energy landscape perspective, these methods usually assume that a large regime shift occurs when a local minimum in which the system’s state is currently sitting is lost via the diminishment of the energy barrier between the currently visited basin and a neighboring basin; see Fig 1 in [103], Figs 2 and 3 in [104], and Box 1 in [106] for schematics. Then, the regime shift corresponds to the movement of the system’s state from the current basin to the neighboring basin. Cross-fertilization of ELA and early warning signal research may be interesting.
@@ Acknowledgements
We acknowledge Pitambar Khanra for technical information.